OpenStudy (anonymous):

Solve for the variables. 2x + 24 = 3y x = 18 - 6y Question 3 options: a) x = -6; y = 4 b) x = 3; y = 10 c) Always True d) None of these In a question like this, what does "Always true" mean?

OpenStudy (mertsj):

Replace x in the first equation with 18-6y from the second equation.

jimthompson5910 (jim_thompson5910):

2x + 24 = 3y 2(18 - 6y) + 24 = 3y ... plug in x = 18 - 6y 36 - 12y + 24 = 3y -12y + 60 = 3y keep going to solve for y

jimthompson5910 (jim_thompson5910):

after you found the value of y, plug it into x = 18 - 6y to find the value of x

OpenStudy (mertsj):

"Always true" means no matter what you replace x and y with, you get true sentences.

OpenStudy (anonymous):

in a question like this, it is when the two lines are on top of each other, i.e, the exact same equation. the two lines presented to you actually intersect at only one point. which is either a or b, i didn't solve. [you can prove this to yourself by drawing them out, easily done by isolating y: first being y= (2x + 24)/3]